OpenStudy (anonymous):

Subtract the following radicals: 6√8 - 4√2 A. 6 B. 6√2 C. 8 D. 8√2

OpenStudy (anonymous):

we have : 6√8 = > 8=2^3 =>√8 = 2√2 .so 6√8=8√2 we have : 8√2 -4√2=4√2 Got it ?

OpenStudy (anonymous):

is \[4\sqrt{2}\] the answer?

OpenStudy (anonymous):

@sione

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

did you see the answers above?

OpenStudy (anonymous):

again : we have : 6√8 - 4√2=6√8 - 2√8=4√8 and can say : 4√8 =6√2 I had a small fall ! The answer is 6√2 .